OpenStudy (anonymous):

The value of 137−−−√ lies between which two consecutive integers?

OpenStudy (perl):

sqrt(137) ?

OpenStudy (anonymous):

i think i know

OpenStudy (anonymous):

what is it?

OpenStudy (cwrw238):

Hint: what is the next perfect square above 137?

OpenStudy (anonymous):

i dont't know:/

OpenStudy (anonymous):

Find the slope of the line passing through the points (3, 8) and (-2, 5). 3 3/5 13 help me please

OpenStudy (anonymous):

u have to figure out what the square above 137 is

OpenStudy (cwrw238):

thats 12 * 12 = ?

OpenStudy (anonymous):

yes i belive so

OpenStudy (anonymous):

whats the answer please

OpenStudy (anonymous):

12*12?

OpenStudy (anonymous):

what?

OpenStudy (cwrw238):

12 * 12 = 144

OpenStudy (anonymous):

i really dont know i am really sorry i thiught it would be 12 star 12?

OpenStudy (cwrw238):

11 * 11 = ?

OpenStudy (anonymous):

i think so but im still trying to figure it out

OpenStudy (anonymous):

The value of 137−−−√ lies between which two consecutive integers?

OpenStudy (anonymous):

i think i got it

OpenStudy (anonymous):

The value of 137 lies between which two consecutive integers?

OpenStudy (anonymous):

what is it then?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

okay

OpenStudy (cwrw238):

@taylor34 - what is 11 times 11 ?

OpenStudy (anonymous):

121

OpenStudy (anonymous):

121 is the ansewer

OpenStudy (cwrw238):

right - and 137 lies between the 2 squares 121 and 144

OpenStudy (anonymous):

no it in't there has to be 2 different numbers for the answer

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

the answer was 11 and 12

OpenStudy (cwrw238):

so the square root of 137 must lie between 11 and 12

OpenStudy (anonymous):

The value of 80 lies between which two consecutive integers

OpenStudy (anonymous):

The value of 80−−√ lies between which two consecutive integers?

OpenStudy (anonymous):

im coming up with an ansewer i believe that taylorg34 is correct

OpenStudy (anonymous):

YES ITS RIGHT!!!!!!!!!!!!

OpenStudy (cwrw238):

yes 11 and 12 are the correct answers - but do you both see why they are correct?